OpenStudy (anonymous):

How do you find the inverse of f(x)= 1+ square root of 1-x2

OpenStudy (anonymous):

\[x=1+\sqrt{1-y^2}\] Solve for y.

OpenStudy (anonymous):

the equation is \[f(x)=1+\sqrt{1-x ^{2}}\]

OpenStudy (anonymous):

hupeorekef

OpenStudy (anonymous):

Yes but in inverse, the x and y switch

OpenStudy (anonymous):

ookay...

OpenStudy (anonymous):

\[x^{2}-2x=-y ^{2}\]

OpenStudy (anonymous):

this is what i got so far.

OpenStudy (anonymous):

how do you solve for y?

OpenStudy (anonymous):

How -2x?

OpenStudy (anonymous):

\[(x-1)^{2}\]

OpenStudy (anonymous):

Ok, that should be fine then. Now make -y^2 positive by multiplying both sides by -1 so you can take the square root.

OpenStudy (anonymous):

And remember that square-roots have 2 solutions: one positive and one negative.

OpenStudy (anonymous):

\[y=\pm \sqrt{-x ^{2}+2x}\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Should all look like this up to that point: Starting with \[x=1+\sqrt{1-y^2}\] Subtract 1 from both sides, then square both sides \[(x-1)^2=1-y^2\] Subtract 1 from both sides, then multiply both sides by -1 (Could also multiply everything by -1 first and then add 1 to both sides..) \[y^2=-(x-1)^2+1 \rightarrow y=±\sqrt{-(x-1)^2+1}\] You can then simplify the radicand if you like.

OpenStudy (anonymous):

Yep, looks like you got it.

OpenStudy (anonymous):

how come when i plug in like -1 or 0 in there the equation is wrong?

OpenStudy (anonymous):

What do you mean by "wrong?"

OpenStudy (anonymous):

if you plug in -1 for x you will get a negative answer in the square root...

OpenStudy (anonymous):

Then -1 is not in the domain of x over the real numbers.

OpenStudy (anonymous):

in this question they gave me the range from \[-1\le(x) \le0\]

OpenStudy (anonymous):

In your original function, any x >1 is outside the domain. Establishing the domain of the function and its inverse is part of the process.

OpenStudy (anonymous):

how would i graph the inverse?

OpenStudy (anonymous):

The original function is the top half of a circle centered at (0,1) with a radius of 1. The inverse is the right half of a circle centered at (1,0) with a radius of 1.

OpenStudy (anonymous):

okay thx